Solve for the unknown side in the following right triangle. (side AC is the hypotenuse.) Round your answer to two places, where applicable.Side AB ? Side BC12 Side AC 19
Answer by nerdybill(7384)

Solve for the unknown side in the following right triangle. (side AC is the hypotenuse.) Round your answer to two places, where applicable.Side AB ? Side BC12 Side AC 19
From Pythagorean theorem we know:
(AC)^2 = (AB)^2 + (BC)^2
plugging in what was given:
(19)^2 = (AB)^2 + (12)^2
(19)^2 = (AB)^2 + (12)^2
361 = (AB)^2 + 144
361 - 144 = (AB)^2
217 = (AB)^2
14.73 = (AB)